How many solutions are there to this system? Please provide all steps.
4x - 7y = 15
-8x + 14y = -30

Answer by TimothyLamb(4379)

---
4x - 7y = 15
-8x + 14y = -30
---
copy and paste the above linear system in standard form into this matrix-method solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution by matrix method:
---
This linear system is singular. It has no solution
---
system matrix
4 -7
-8 14
---
determinant of system matrix = 0
---
now by solution by elimination:
---
( 4x - 7y = 15 )2
-8x + 14y = -30
---
8x - 14y = 30
-8x + 14y = -30
---
add the system:
---
0 = 0
---
This linear system is singular. It has no solution
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php